How to Write an Abstract for the Conference on High Performance Scientific Computing

N. H. Dien^{1}, T. D. Phuong^{2}, and \underline{T. H. Thai}^{3}

Abstract: This text is a style example for abstracts of the International Conference on High Performance Scientific Computing (March 10-14, 2003, Hanoi), which must be written in L^{A}TEX-format, not to exceed one A4-page, and transmitted electronically or sent as files in a diskette. You can load the needed style file abs4hpsc.sty and this example file abs4hpsc.tex from our Web-page or ask us to send them to you. It is recommended to copy this L^{A}TEX-file and modify it by your new content.

The title and the list of author’s must be placed between the brackets { } after the command \newtitle or \newauthor, respectively. For long title and long list, use \\ to demand line breaking followed by \medskip to have a suitable skip. The author who will actually give the talk is marked by an underline (\underbar{ }). The upper index after every author’s name made by a number in the brackets { } after the command \no following the name is to indicate the corresponding address in the list at the end. To make this list, you have to use the command \newaddress{ } and the environment itemize with \no{ } in the bracket of \item[ ]. For instance, the list of addresses at the end of this text is written by

\newaddress{
  \begin{itemize}
    \item[\no{1,2}] Department of ... and Scientific Computing \\
      ...
      {\it nhdien@thevinh.ac.vn, tdphuong@thevinh.ac.vn}
    \item[\no{3}] Interdisciplinary Center ... Heidelberg\\
      Im Neuenheimer Feld 368, 69120 Heidelberg, Germany\\
      {\it thai@iwr.uni-heidelberg.de}
  \end{itemize} }

Please use {\it ...} for e-mail address.

The proper content of abstract starts with the command \newabstract. To begin a new paragraph just add a blank row in L^{A}TEX-file without any skip command.

————————————–

1,2 Department of Numerical Analysis and Scientific Computing Institute of Mathematics, National Center for Sciences and Technoloy P. O. Box 631 - Bo Ho, Hanoi, Vietnam nhdien@thevinh.ac.vn, tdphuong@thevinh.ac.vn

3 Interdisciplinary Center for Scientific Computing, University of Heidelberg Im Neuenheimer Feld 368, 69120 Heidelberg, Germany thai@iwr.uni-heidelberg.de